OpenStudy (anonymous):

A ladder 30 feet long leans against a vertical wall, with it's foot 5 feet from the base of the wall. At what distance from the ground does the top of the ladder touch the wall?

jimthompson5910 (jim_thompson5910):

We have this right triangle 

jimthompson5910 (jim_thompson5910):

so use the pythagorean theorem to solve for x a^2 + b^2 = c^2 x^2 + 5^2 = 30^2 x^2 + 25 = 900 ... ... ... x = ??